OpenStudy (anonymous):

how is 3-2root2 = (root2-1)^2

OpenStudy (anonymous):

\[3-2\sqrt{2}=(\sqrt{2}-1)^2\]

OpenStudy (anonymous):

how to know that lhs can also be written in the form of rhs

OpenStudy (butterflydreamer):